OpenStudy (anonymous):

how do I graph y=(x-3)^2

OpenStudy (anonymous):

expand it ,and then find the corresponding values of Y for different values of X ,and then graph it using appropriate scale ,taking y along vertical axis and x on horizental

OpenStudy (anonymous):

so, then it would be y=(x-3)(x-3) or, y=x^2-6x+9...?

OpenStudy (anonymous):

if using values for x of: -1,0,1 then co-ordinates would be the corresponding pairs of: (-1,4), (0,9) and (1,4) ?

OpenStudy (anonymous):

take values of x as 0,1,2,3,4 then , the answer u get is the value of y then sketch it and give me a mdal :p

OpenStudy (anonymous):

Make a table, when x=0, y=(0-3)^2 x=1, y=(1-3)^2 Go as far as you want. Usually, people would do 0, then two positive x, and two negative x